OpenStudy (anonymous):

Suppose that I have an infinite collection of open intervals \(O_n\) whose lengths are \(1/2^{n-1}\), where \(n\in\mathbb{N}\). Would then\[O=\bigcup_{n=1}^{\infty}O_{n}\]have a length of \(2\)? If so, does this imply that\[O^{c}=\bigcap_{n=1}^{\infty}O_{n}^{c}\]has infinite length?

OpenStudy (anonymous):

i say that \(O\) would have length \(2\) because \(1+1/2+1/2^2+...=2\).

OpenStudy (anonymous):

confused. you say the lengths are each \[\frac{1}{2^{n-1}}\] but who says they are disjoint?

OpenStudy (anonymous):

hmm very good point

OpenStudy (anonymous):

suppose the sets are the intevals \[(\frac{1}{2^{n-1}}, \frac{1}{2^{n-1}})\] then since they are nested the union would be the first (largest) one

OpenStudy (anonymous):

i forgot to add that the open intervals are centered at unique rational numbers, but your point still stands.

OpenStudy (anonymous):

i meant of course \[(-\frac{1}{2^{n-1}}, \frac{1}{2^{n-1}})\]

OpenStudy (anonymous):

i could argue that the greatest length the union of these intervals can attain is of 2, though.

OpenStudy (anonymous):

actually i think it would be one

OpenStudy (anonymous):

\[(-\frac{1}{2},\frac{1}{2})\] has length 1

OpenStudy (anonymous):

that'd be the smallest (given they're all nested, like you described).

OpenStudy (anonymous):

i think that would be the largest

OpenStudy (anonymous):

if they're disjoint, then the total length would be 2.

OpenStudy (anonymous):

wait... hmm...

OpenStudy (anonymous):

yes if they were disjoint you could add them up as you did in a geometric series get 2

OpenStudy (anonymous):

gotta run, good luck

OpenStudy (anonymous):

kk thx